OpenStudy (mama2dylan):

You have a coupon good for $5 off the price, p, of any large pizza. You also get a 15% discount on any pizza if you show your student ID. How much more would you pay for a large pizza if the cashier applies the coupon first, rather than the 15% discount first? The answer is 75 cents please show how steps using a composite function.

OpenStudy (funnyguy2):

how much money does it start out with

OpenStudy (mama2dylan):

That's the thing, it doesn't say!

OpenStudy (sleepyjess):

I can't view that link. Could you take a screenshot?

OpenStudy (mama2dylan):

Sure @sleepyjess hold on.

OpenStudy (mama2dylan):

This is a 2 page example from the book of what they are looking for using different figures.

Directrix (directrix):

Let p = cost of pizza. Coupon cost of pizza = p - 5 Let c(p) = p - 5 Discount cost of pizza = p - .15p = .85p Let d(p) = .85 p

Directrix (directrix):

C (D(p) = C (.85p ) = .85p - 5

Directrix (directrix):

D ( C(p)) = D (p - 5) = .85 * (p - 5) = .85p - 4.25

Directrix (directrix):

Compare: C (D(p) =85p - 5 with D ( C(p) = .85p - 4.25

Directrix (directrix):

.85p - 5 versus .85p - 4.25 Differ by 5 - 4.25 = $.75 = 75 cents @mama2dylan

Directrix (directrix):

Coupon followed by 15% discount saves 75 cents.